OpenStudy (anonymous):

i started this equation, can anyone tell me if im doing it right.. im supposed to have a minimum of 7 lines 5(2x - 7) = 33 - 3(6x + 4) 10x - 7 = 33 + 18x + 12

OpenStudy (anonymous):

I believe your on the correct track I'm about to start trying to solve it myself

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) 10x-35 = 33-18x-12

OpenStudy (anonymous):

5(2x-7) => (5*2x)-(5*7) -> 10x - 35

OpenStudy (anonymous):